OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (wach):

We might first try and put this in a form we're familiar with. Do you know point intercept form?

OpenStudy (anonymous):

no

OpenStudy (wach):

Like y = mx + b where m is the slope

OpenStudy (wach):

2x - 5y = -10 So we'll solve for y.

OpenStudy (anonymous):

alright wouldnt it be 2

OpenStudy (wach):

2x - 5y = -10 subtract 2x -5y = -2x - 10 divide by -5 y = -2x/-5 - 10/-5 y = 2/5x +2

OpenStudy (wach):

The place where the line crosses the y intercept would be 2, but we want m or slope

OpenStudy (wach):

Compare y = 2/5x +2 y = mx +b which number would be m? That's the slope.

OpenStudy (wach):

So, the answer is 2/5, right?

OpenStudy (anonymous):

ya lol thanx for help it helped alot

OpenStudy (wach):

Cool :)